OpenStudy (anonymous):

two flasks are connected with a stopwingspan. flask #1 has a volume of 2.5L and contains oxygen gas at a pressure of .7atm. Flask #2 has a colume of 3.8L and contains hydrogen gas at a pressure of 1.25atm. when the stopwingspan between the two flasks is opened and the gases mix, what will the resulting pressure of the gas mixture be?

OpenStudy (anonymous):

 Individually each gas is ideal and follows the characters of the ideal gas law. flask#1 PV=nRT n = PV/RT for flask # 2 PV=nRT n = PV/RT ------------------- When the two flasks are connected the volume add are combined and the number of moles of each gas are combined. just calculate the new pressure. P =nRT/V